OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5. A. y = 6x + 5 B. y = -5/6 x + 5 C. y = -6x + 5 D. y = 5/6 x + 5

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (secret-ninja):

I like to break these down. So slope intercept form is: y = mx + b And m is the slope and b is the y-intercept They give us the y intercept, so we can start filling in the equation: y = mx + 5 So now we substitute in values for x and y to solve for m: We know the x-intercept is 6, so that's the point (6,0), where x=6 and y=0, so let's substitute those into our equation so far: 0 = m(6) + 5 -5 = 6m m = -5/6 So the equation is: y = (-5/6)x + 5 A.K.A. B. :)

OpenStudy (anonymous):

wow that made sense to thank you

OpenStudy (secret-ninja):

No problem. :)